# Anomalous Coherence Length in Superconductors with Quantum Metric

### v5

@@ II.2 Topologically trivial flat-band model
To support the analytical results mentioned above, we employ the mean-field theory on a microscopic model, which features exactly flat bands without dispersion Hofmann et al. (2022, 2023b); Mao and Chowdhury (2023). The normal state Hamiltonian $h_{s}(\bm{k})$ for electrons with spin index $s$ reads
@@
Figure 4: For topological flat-band model: a A two-orbital square lattice with short- and long-range hoppings, b the electronic band structure, c the quantum metric distribution of the lower flat band, and d coherence length $\xi$ v.s. pairing gap $\Delta$. In a, the inter-orbital nearest hopping, intra-orbital next-nearest-neighbor and fifth-nearest-neighbor hoppings are labeled. In b, the lower band (purple) has nearly zero bandwidth with the Chern number $C=2$. In d, the coherence length is extracted from the pair correlation function and it is bounded by a Chern number, which is guided by the dashed green line.
@@ II.3 Topological bound of the coherence length
In the previous subsection, we have demonstrated how the quantum metric gives a lower bound for the superconducting coherence length. We now consider a system which possesses topological flat bands. As pointed out previously Roy (2014); Marzari and Vanderbilt (1997); Peotta and Törmä (2015); Provost and Vallee (1980b), the quantum metric has a lower bound which is proportional to the Chern number. Therefore, we expect that there is a finite quantum metric length which serves as the lower bound of the superconducting coherence length for a superconductor with nontrivial spin Chern numbers.
Specifically, the quantum geometric tensor is a positive semidefinite matrix, and in two spatial dimensions, we have the inequality $\sqrt{\mathrm{det}\mathcal{G}(\bm{k})}\geq|\mathcal{F}_{xy}(\bm{k})|/2$ Roy (2014), which implies that a topological band must necessarily possess a finite quantum metric. According to Eq. (2), this indicates that there is a lower bound on the coherence length $\xi$ which is determined by the topology of the band such that
@@
Figure 5: A list of the moiré band structures, quantum metric and geometric contributions of the coherence length $\ell_{\mathrm{qm}}$ for MATBG, MATTG and TDBG. For both MATBG and MATTG, the quantum metric is plotted for the highest valence band, and it exhibits divergence near the $K$ points. In TDBG, an electric field potential of $V=40$ meV is applied, leading to flat bands near charge neutrality with Chern number $C=\pm 2$. The quantum metric is plotted for the lowest conduction band. In evaluating $\ell_{\mathrm{qm}}$, we ignore the band dispersion.
@@ II.4 Application to Moiré materials
The graphene-based moiré systems provide the versatile platforms to explore the exotic phenomena related to the flat bands Ledwith et al. (2020); Da Liao et al. (2021); Song and Bernevig (2022); Chou and Das Sarma (2023); Herzog-Arbeitman et al. (2024); Hu et al. (2023a); Zhang et al. (2023); Hu et al. (2023b); Kolář et al. (2023); Kwan et al. (2021). In the superconducting graphene-based moiré family Cao et al. (2018); Yankowitz et al. (2019); Oh et al. (2021); Arora et al. (2020), the quantum metric effect is indeed very crucial. Particularly, the quantum metric plays a significant role in determining the coherence length in magic angle twisted bilayer graphene (MATBG) with twisted angle $\theta\approx 1.08^{\circ}$. To provide a qualitative estimation of the impact of the quantum metric, we employ the Bistritzer-MacDonald model to elucidate the significance of the quantum metric in the context of graphene-based moiré materials Bistritzer and MacDonald (2011). We also assume the presence of an $s$-wave superconducting phase. As shown in Fig. 5, the quantum metric length $\ell_{\mathrm{qm}}=1.2L_{M}\approx 13$ nm. Here, $L_{M}\approx a_{0}/\theta$ represents the moiré lattice constant. By employing the self-consistent mean-field study (in the Supplementary Note IV), we calculate the total coherence length using Eq. (8) to take into account the band dispersion. Using the interaction strength $U=0.6$ meV, which gives $T_{c}\approx 1.7K$, we obtain a conventional contribution of $\sim 3$nm and $\ell_{\mathrm{qm}}\sim 13$ nm at $\theta=1.08^{\circ}$. Therefore, the total superconducting coherence length given by Eq. (2), is indeed dominated by the quantum metric contribution.
A large family of moiré systems exhibit superconductivity, such as magic-angle twisted trilayer graphene (MATTG) Park et al. (2021, 2022) and twisted double-bilayer graphene (TDBG) Liu et al. (2020). Similar to MATBG, the quantum metric effects cannot be neglected, as shown in Fig. 5. For MATTG, $\ell_{\mathrm{qm}}=1.2L_{M}$, and for TDBG, $\ell_{\mathrm{qm}}=0.5L_{M}$. The calculations of $\ell_{\mathrm{qm}}$ in Fig. 5 are made by averaging the quantum metric over the moiré Brillouin zone without considering the quasiparticle energy in Eq. (11). Notably, the flat band in TDBG carries a non-zero valley Chern number $C=2$, leading to a topology-bound coherence length, as discussed previously. We focus on the quantum metric within a single band, while the generalization to multiple nearly-degenerate flat bands consists of replacing the one-band quantum metric with the non-abelian quantum metric Mera and Mitscherling (2022); Herzog-Arbeitman et al. (2022b). The quantum metric length calculated for the moiré systems is a qualitative estimation because of the limitations of the continuum model and the simple $s$-wave pairing assumption. It will be an open question of the role that quantum metric plays in unconventional superconductors for moiré systems.
III Method
@@
IV Conclusion
Acknowledgments. We thank Adrian Po for illuminating discussions. K.T.L. acknowledges the support of the Ministry of Science and Technology, China, and the Hong Kong Research Grants Council through Grants No. 2020YFA0309600, No. RFS2021-6S03, No. C6025-19G, No. AoE/P-701/20, No. 16310520, No. 16310219 and No. 16307622.
Supplementary Note for
“Anomalous Coherence Length in Superconductors with Quantum Metric”
Jin-Xin Hu,1 Shuai A. Chen,1 K. T. Law1
1Department of Physics, Hong Kong University of Science and Technology, Clear Water Water Bay, Hong Kong, China
I Mean-field theory
@@
I.1 Mean-field theory for a multi-band superconductor
@@
$H_{0}$ $=\sum_{ij}\sum_{\alpha\beta\sigma}h_{ij\alpha\beta}^{\sigma}a_{i\alpha\sigma}^{\dagger}a_{j\beta\sigma},$ (S1)
$H_{\mathrm{int}}$ $=-\sum_{i\alpha}Ua_{i\alpha\uparrow}^{\dagger}a_{i\alpha\downarrow}^{\dagger}a_{i\alpha\downarrow}a_{i\alpha\uparrow},$ (S2)
@@
$\Delta_{\alpha}=-U\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}\rangle.$ (S3)
@@
$H_{\mathrm{mf}}$ $=\sum_{ij}\sum_{\alpha\beta\sigma}h_{ij\alpha\beta}^{\sigma}a_{i\alpha\sigma}^{\dagger}a_{j\beta\sigma}+\sum_{i\alpha}\left[a_{i\alpha\uparrow}^{\dagger}a_{i\alpha\downarrow}^{\dagger}\Delta_{\alpha}+a_{i\alpha\downarrow}a_{i\alpha\uparrow}\Delta_{\alpha}^{*}\right]$ (S4)
$=\sum_{\bm{k}}\sum_{\alpha\beta\sigma}h_{\alpha\beta}^{\sigma}(\bm{k})a_{\alpha\sigma}^{\dagger}(\bm{k})a_{\beta\sigma}(\bm{k})+\sum_{\bm{k}\alpha}\left[a_{\alpha\uparrow}^{\dagger}(\bm{k})a_{\alpha\downarrow}^{\dagger}(-\bm{k})\Delta_{\alpha}+a_{\alpha\downarrow}(\bm{k})a_{\alpha\uparrow}(-\bm{k})\Delta_{\alpha}^{*}\right],$ (S5)
@@
$a_{\alpha\sigma}^{\dagger}(\bm{k})$ $=\frac{1}{\sqrt{N}}\sum_{i}e^{i(\bm{r}_{i}+\bm{\delta}_{\alpha})\cdot\bm{k}}a_{i\alpha\sigma}^{\dagger}$ (S6)
$a_{i\alpha\sigma}^{\dagger}$ $=\frac{1}{\sqrt{N}}\sum_{\bm{k}}e^{-i(\bm{r}_{i}+\bm{\delta}_{\alpha})\cdot\bm{k}}a_{\alpha\sigma}^{\dagger}(\bm{k}).$ (S7)
@@
$h_{\alpha\beta}^{\sigma}(\bm{k})=\sum_{ij}h_{ij\alpha\beta}^{\sigma}e^{-i(\bm{r}_{i}+\bm{\delta}_{\alpha}-\bm{r}_{j}-\bm{\delta}_{\beta})\cdot\bm{k}}.$ (S8)
@@
$H_{\mathrm{mf}}=\sum_{\bm{k}}\psi_{\bm{k}}^{\dagger}(\hat{h}\otimes\tau_{z}+\mathrm{Re}\hat{\Delta}\otimes\tau_{x}+\mathrm{Im}\hat{\Delta}\otimes\tau_{y})\psi_{\bm{k}}+\sum_{\bm{k}}\sum_{\alpha\beta}\left[h_{\alpha\beta}(\bm{k})-\mu\delta_{\alpha\beta}\right]-\frac{1}{U}\sum_{\alpha}|\Delta_{\alpha}|^{2},$ (S9)
@@
$G(i\omega_{n},\bm{k})=\frac{1}{i\omega_{n}+\hat{h}\otimes\tau_{z}+\mathrm{Re}\hat{\Delta}\otimes\tau_{x}+\mathrm{Im}\hat{\Delta}\otimes\tau_{y}}.$ (S10)
@@
$\hat{\Delta}_{\alpha}=\frac{T}{N}\sum_{n}\sum_{\sigma\sigma^{\prime}}\sum_{\bm{k}}\hat{\hat{G}}_{\alpha\sigma,\alpha\sigma^{\prime}}(i\omega_{n},\bm{k})(\tau_{x})_{\sigma^{\prime}\sigma},$ (S11)
@@
$2\nu=1+\frac{T}{N}\sum_{n}\sum_{\sigma\sigma^{\prime}}\sum_{\alpha}\sum_{\bm{k}}\hat{G}_{\alpha\sigma,\alpha\sigma^{\prime}}(i\omega_{n},\bm{k})(\tau_{z})_{\sigma\sigma^{\prime}}.$ (S12)
@@
$C(\bm{r},\bm{r}^{\prime})$ $=\sum_{\alpha}\langle a_{\alpha\uparrow}(\bm{r})a_{\alpha\downarrow}(\bm{r})a_{\alpha\downarrow}^{\dagger}(\bm{r}^{\prime})a_{\alpha\uparrow}^{\dagger}(\bm{r}^{\prime})\rangle$
$=\frac{T}{N}\sum_{n}\sum_{\bm{q}}e^{-i\bm{q}\cdot(\bm{r}-\bm{r}^{\prime})}\sum_{\alpha\bm{k}}[G_{0}^{\alpha\alpha}(i\omega_{n},\bm{k}+\bm{q})G_{0}^{\alpha\alpha}(-i\omega_{n},-\bm{k})+F_{0}^{\alpha\alpha}(i\omega_{n},\bm{k}+\bm{q})F_{0}^{\dagger\alpha\alpha}(-i\omega_{n},-\bm{k})],$ (S13)
with Gor’kov’s normal and anomalous Green’s functions $G_{0}$ and $F_{0}$. This formula can be employed to evaluate the coherence length for several lattice model, such as sawtooth and Lieb lattices.
I.2 Projected mean-field Hamiltonian and Green function
@@
$a_{i\alpha\sigma}^{\dagger}\rightarrow\frac{1}{\sqrt{N}}\sum_{\bm{k}}e^{-i(\bm{r}_{i}+\bm{\delta}_{\alpha})\cdot\bm{k}}u_{\alpha\sigma}(\bm{k})c_{\sigma}^{\dagger}(\bm{k}),$ (S14)
@@
$H_{0}=\sum_{\bm{k}\sigma}\epsilon(\bm{k})c_{\bm{k}\sigma}^{\dagger}c_{\bm{k}\sigma}+\sum_{\bm{k}}\left[c_{\uparrow}^{\dagger}(\bm{k})c_{\downarrow}^{\dagger}(-\bm{k})\Delta_{0}+h.c.\right],$ (S15)
@@
$\Delta_{0}=\frac{1}{N}\sum_{\alpha}\sum_{\bm{k}}\Delta_{\alpha}u_{\alpha}(\bm{k})u_{\alpha}^{*}(\bm{k}).$ (S16)
@@
$\Delta=-\frac{U}{N}\sum_{\bm{k}}\langle c_{\downarrow}(-\bm{k})c_{\uparrow}(\bm{k})\rangle,$ (S17)
@@
$=\sum_{\bm{k}}\psi_{\bm{k}}^{\dagger}[(\epsilon(\bm{k})-\mu)\tau_{z}+\Delta\tau_{x}]\psi_{\bm{k}}+\sum_{\bm{k}}(\epsilon(\bm{k})-\mu)+\frac{1}{U}\Delta^{2},$ (S18)
@@
$G(\bm{r}\tau;\bm{r}^{\prime}\tau^{\prime})=\begin{bmatrix}\langle\hat{T}[c_{\uparrow}(\bm{r},\tau)c_{\uparrow}^{\dagger}(\bm{r}^{\prime},\tau^{\prime})]\rangle&\langle\hat{T}[c_{\uparrow}(\bm{r},\tau)c_{\downarrow}(\bm{r}^{\prime},\tau^{\prime})]\rangle\\ \langle\hat{T}[c_{\downarrow}^{\dagger}(\bm{r},\tau)c_{\uparrow}^{\dagger}(\bm{r}^{\prime},\tau^{\prime})]\rangle&\langle\hat{T}[c_{\downarrow}^{\dagger}(\bm{r},\tau)c_{\downarrow}(\bm{r}^{\prime},\tau^{\prime})]\rangle\end{bmatrix},$ (S19)
@@
$G(i\omega_{n},\bm{k})=\frac{-i\omega_{n}\tau_{0}-(\epsilon(\bm{k})-\mu)\tau_{z}-\Delta\tau_{x}}{\omega_{n}^{2}+(\epsilon(\bm{k})-\mu)^{2}+\Delta^{2}},$ (S20)
where $\omega_{n}=(2n+1)\pi T$ is the Matsubara frequency and $\tau_{0}$ is a $2\times 2$ identity matrix. With the Gor’kov’s Green function in Eq. (S20), we can determine the order parameters by
$\Delta=-\frac{U}{2N}T\sum_{\bm{k},n}\mathrm{Tr}G(i\omega_{n},\bm{k})\tau_{x}=\frac{U}{N}\sum_{\bm{k},n}\frac{\Delta}{\omega_{n}^{2}+(\epsilon(\bm{k})-\mu)^{2}+\Delta^{2}},$ (S21)
@@
$1=\frac{U}{N}T\sum_{\bm{k},n}\frac{1}{\omega_{n}^{2}+(\epsilon(\bm{k})-\mu)^{2}+\Delta^{2}}.$ (S22)
@@
$2\nu=1+\frac{T}{N}\sum_{\bm{k},n}\mathrm{Tr}G(i\omega_{n},\bm{k})\tau_{z}.$ (S23)
with $\nu$ being the filling factor. Combine Eqs. (S22) and (S23) and we can obtain the superconducting order parameter.
I.3 Interaction renormalized chemical potential $\mu$
@@
$\nu$ $=\frac{1}{2}\left(1-\text{$\int\frac{d^{2}k}{(2\pi)^{2}}$}\frac{\epsilon(\bm{k})-\mu_{0}}{|\epsilon(\bm{k})-\mu_{0}|}\tanh\frac{\beta|\epsilon(\bm{k})-\mu_{0}|}{2}\right),$ (S24)
$\nu$ $=\frac{1}{2}\left(1-\int\frac{d^{2}k}{(2\pi)^{2}}\frac{\epsilon(\bm{k})-\mu}{\varepsilon(\bm{k})}\tanh\frac{\beta\varepsilon(\bm{k})}{2}\right).$ (S25)
@@
$=\frac{m}{2\pi}\left(\sqrt{\Delta^{2}}-\sqrt{\Delta^{2}+\mu^{2}}\right).$ (S26)
@@
$\mu_{0}=\Delta-\sqrt{\Delta^{2}+\mu^{2}}=\frac{\mu^{2}}{\Delta+\sqrt{\Delta^{2}+\mu^{2}}},$ (S27)
@@
II Quantum geometry and coherence length
II.1 Proof of the independence of coherence length from orbital positions

### v6

@@ II.2 Topologically trivial flat-band model
To support the analytical results mentioned above, we employ the mean-field theory on a microscopic model, which features exactly flat bands without dispersion Hofmann et al. (2022, 2023). The normal state Hamiltonian $h_{s}(\bm{k})$ for electrons with spin index $s$ reads
@@
Figure 4: Quantum metric and coherence length for topological flat-band model. a A two-orbital square lattice with short- and long-range hoppings, b the electronic band structure, c the quantum metric distribution of the lower flat band, and d coherence length $\xi$ v.s. pairing gap $\Delta$. In a, the inter-orbital nearest hopping, intra-orbital next-nearest-neighbor and fifth-nearest-neighbor hoppings are labeled. In b, the lower band (purple) has nearly zero bandwidth with the Chern number $C=2$. In d, the coherence length is extracted from the pair correlation function and it is bounded by a Chern number, which is guided by the dashed green line.
@@ II.3 Topological bound of the coherence length
In the previous subsection, we have demonstrated how the quantum metric gives a lower bound for the superconducting coherence length. We now consider a system which possesses topological flat bands. As pointed out previously Roy (2014); Marzari and Vanderbilt (1997), the quantum metric has a lower bound which is proportional to the Chern number. Therefore, we expect that there is a finite quantum metric length which serves as the lower bound of the superconducting coherence length for a superconductor with nontrivial spin Chern numbers.
Specifically, the quantum geometric tensor is a positive semidefinite matrix, and in two spatial dimensions, we have the inequality $\sqrt{\mathrm{det}\mathcal{G}(\bm{k})}\geq|\mathcal{F}_{xy}(\bm{k})|/2$, which implies that a topological band must necessarily possess a finite quantum metric. According to Eq. (2), this indicates that there is a lower bound on the coherence length $\xi$ which is determined by the topology of the band such that
@@
Figure 5: A list of the moiré band structures, quantum metric and geometric contributions of the coherence length $\ell_{\mathrm{qm}}$ for magic angle twisted bilayer graphene (MATBG), magic angle twisted trilayer graphene (MATTG) and twisted double bilayer graphene (TDBG). For both MATBG and MATTG, the quantum metric is plotted for the highest valence band, and it exhibits divergence near the $K$ points. In TDBG, an electric field potential of $V=40$ meV is applied, leading to flat bands near charge neutrality with Chern number $C=\pm 2$. The quantum metric is plotted for the lowest conduction band. In evaluating $\ell_{\mathrm{qm}}$, we ignore the band dispersion.
@@ II.4 Application to Moiré materials
The graphene-based moiré systems provide the versatile platforms to explore the exotic phenomena related to the flat bands Ledwith et al. (2020); Da Liao et al. (2021); Song and Bernevig (2022); Chou and Das Sarma (2023); Herzog-Arbeitman et al. (2024); Hu et al. (2023a); Zhang et al. (2023); Hu et al. (2023b); Kolář et al. (2023); Kwan et al. (2021). In the superconducting graphene-based moiré family Park et al. (2022), the quantum metric effect is indeed very crucial. Particularly, the quantum metric plays a significant role in determining the coherence length in magic angle twisted bilayer graphene (MATBG) with twisted angle $\theta\approx 1.08^{\circ}$. To provide a qualitative estimation of the impact of the quantum metric, we employ the Bistritzer-MacDonald model to elucidate the significance of the quantum metric in the context of graphene-based moiré materials Bistritzer and MacDonald (2011). We also assume the presence of an $s$-wave superconducting phase. As shown in Fig. 5, the quantum metric length $\ell_{\mathrm{qm}}=1.2L_{M}\approx 13$ nm. Here, $L_{M}\approx a_{0}/\theta$ represents the moiré lattice constant. By employing the self-consistent mean-field study (in the Supplementary Note 4), we calculate the total coherence length using Eq. (8) to take into account the band dispersion. Using the interaction strength $U=0.6$ meV, which gives $T_{c}\approx 1.7K$, we obtain a conventional contribution of $\sim 3$nm and $\ell_{\mathrm{qm}}\sim 13$ nm at $\theta=1.08^{\circ}$. Therefore, the total superconducting coherence length given by Eq. (2), is indeed dominated by the quantum metric contribution.
A large family of moiré systems exhibit superconductivity, such as magic-angle twisted trilayer graphene (MATTG) Park et al. (2021) and twisted double-bilayer graphene (TDBG) Liu et al. (2020). Similar to MATBG, the quantum metric effects cannot be neglected, as shown in Fig. 5. For MATTG, $\ell_{\mathrm{qm}}=1.2L_{M}$, and for TDBG, $\ell_{\mathrm{qm}}=0.5L_{M}$. The calculations of $\ell_{\mathrm{qm}}$ in Fig. 5 are made by averaging the quantum metric over the moiré Brillouin zone without considering the quasiparticle energy in Eq. (11). Notably, the flat band in TDBG carries a non-zero valley Chern number $C=2$, leading to a topology-bound coherence length, as discussed previously. We focus on the quantum metric within a single band, while the generalization to multiple nearly-degenerate flat bands consists of replacing the one-band quantum metric with the non-abelian quantum metric Mera and Mitscherling (2022); Herzog-Arbeitman et al. (2022b). The quantum metric length calculated for the moiré systems is a qualitative estimation because of the limitations of the continuum model and the simple $s$-wave pairing assumption. It will be an open question of the role that quantum metric plays in unconventional superconductivity for moiré systems.
III Conclusion
IV Methods
@@
V Data Availability
The data generated from our codes that support the findings of this study are available from the corresponding author upon reasonable request.
VI Author Contributions
K.T.L. and S.C. conceived the project. J.-X.H. performed the major part of the calculations and analysis. J.-X.H., S.C. and K.T.L. wrote the manuscript with contributions from all authors. All authors are involved in the discussions.
VII Acknowledgement
We thank Adrian Po for illuminating discussions. K.T.L. acknowledges the support of the Ministry of Science and Technology, China, and the Hong Kong Research Grants Council through Grants No. 2020YFA0309600, No. RFS2021-6S03, No. C6025-19G, No. AoE/P-701/20, No. 16310520, No. 16309223 and No. 16307622.
VIII Competing Interests
The authors declare no competing interests.
Supplementary Material: Anomalous Coherence Length in Superconductors with Quantum Metric
Supplementary Note 1: Mean-field theory
@@ VIII Competing Interests
Mean-field theory for a multi-band superconductor
@@
$H_{0}$ $=\sum_{ij}\sum_{\alpha\beta\sigma}h_{ij\alpha\beta}^{\sigma}a_{i\alpha\sigma}^{\dagger}a_{j\beta\sigma},$ (1)
$H_{\mathrm{int}}$ $=-\sum_{i\alpha}Ua_{i\alpha\uparrow}^{\dagger}a_{i\alpha\downarrow}^{\dagger}a_{i\alpha\downarrow}a_{i\alpha\uparrow},$ (2)
@@
$\Delta_{\alpha}=-U\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}\rangle.$ (3)
@@
$H_{\mathrm{mf}}$ $=\sum_{ij}\sum_{\alpha\beta\sigma}h_{ij\alpha\beta}^{\sigma}a_{i\alpha\sigma}^{\dagger}a_{j\beta\sigma}+\sum_{i\alpha}\left[a_{i\alpha\uparrow}^{\dagger}a_{i\alpha\downarrow}^{\dagger}\Delta_{\alpha}+a_{i\alpha\downarrow}a_{i\alpha\uparrow}\Delta_{\alpha}^{*}\right]$ (4)
$=\sum_{\bm{k}}\sum_{\alpha\beta\sigma}h_{\alpha\beta}^{\sigma}(\bm{k})a_{\alpha\sigma}^{\dagger}(\bm{k})a_{\beta\sigma}(\bm{k})+\sum_{\bm{k}\alpha}\left[a_{\alpha\uparrow}^{\dagger}(\bm{k})a_{\alpha\downarrow}^{\dagger}(-\bm{k})\Delta_{\alpha}+a_{\alpha\downarrow}(\bm{k})a_{\alpha\uparrow}(-\bm{k})\Delta_{\alpha}^{*}\right],$ (5)
@@
$a_{\alpha\sigma}^{\dagger}(\bm{k})$ $=\frac{1}{\sqrt{N}}\sum_{i}e^{i(\bm{r}_{i}+\bm{\delta}_{\alpha})\cdot\bm{k}}a_{i\alpha\sigma}^{\dagger}$ (6)
$a_{i\alpha\sigma}^{\dagger}$ $=\frac{1}{\sqrt{N}}\sum_{\bm{k}}e^{-i(\bm{r}_{i}+\bm{\delta}_{\alpha})\cdot\bm{k}}a_{\alpha\sigma}^{\dagger}(\bm{k}).$ (7)
@@
$h_{\alpha\beta}^{\sigma}(\bm{k})=\sum_{ij}h_{ij\alpha\beta}^{\sigma}e^{-i(\bm{r}_{i}+\bm{\delta}_{\alpha}-\bm{r}_{j}-\bm{\delta}_{\beta})\cdot\bm{k}}.$ (8)
@@
$H_{\mathrm{mf}}=\sum_{\bm{k}}\psi_{\bm{k}}^{\dagger}(\hat{h}\otimes\tau_{z}+\mathrm{Re}\hat{\Delta}\otimes\tau_{x}+\mathrm{Im}\hat{\Delta}\otimes\tau_{y})\psi_{\bm{k}}+\sum_{\bm{k}}\sum_{\alpha\beta}\left[h_{\alpha\beta}(\bm{k})-\mu\delta_{\alpha\beta}\right]-\frac{1}{U}\sum_{\alpha}|\Delta_{\alpha}|^{2},$ (9)
@@
$G(i\omega_{n},\bm{k})=\frac{1}{i\omega_{n}+\hat{h}\otimes\tau_{z}+\mathrm{Re}\hat{\Delta}\otimes\tau_{x}+\mathrm{Im}\hat{\Delta}\otimes\tau_{y}}.$ (10)
@@
$\hat{\Delta}_{\alpha}=\frac{T}{N}\sum_{n}\sum_{\sigma\sigma^{\prime}}\sum_{\bm{k}}\hat{G}_{\alpha\sigma,\alpha\sigma^{\prime}}(i\omega_{n},\bm{k})(\tau_{x})_{\sigma^{\prime}\sigma},$ (11)
@@
$2\nu=1+\frac{T}{N}\sum_{n}\sum_{\sigma\sigma^{\prime}}\sum_{\alpha}\sum_{\bm{k}}\hat{G}_{\alpha\sigma,\alpha\sigma^{\prime}}(i\omega_{n},\bm{k})(\tau_{z})_{\sigma\sigma^{\prime}}.$ (12)
@@
$C(\bm{r},\bm{r}^{\prime})$ $=\sum_{\alpha}\langle a_{\alpha\uparrow}(\bm{r})a_{\alpha\downarrow}(\bm{r})a_{\alpha\downarrow}^{\dagger}(\bm{r}^{\prime})a_{\alpha\uparrow}^{\dagger}(\bm{r}^{\prime})\rangle_{c}$
$=\frac{T}{N}\sum_{n}\sum_{\bm{q}}e^{-i\bm{q}\cdot(\bm{r}-\bm{r}^{\prime})}\sum_{\alpha\bm{k}}G_{0}^{\alpha\alpha}(i\omega_{n},\bm{k}+\bm{q})G_{0}^{\alpha\alpha}(-i\omega_{n},-\bm{k}),$ (13)
with Gor’kov’s normal Green’s functions $G_{0}$. Here $\langle\mathcal{O}(\mathbf{r})\mathcal{O}(\mathbf{r}^{\prime})\rangle_{c}\equiv\langle\mathcal{O}(\mathbf{r})\mathcal{O}(\mathbf{r}^{\prime})\rangle-\langle\mathcal{O}(\mathbf{r})\rangle\langle\mathcal{O}(\mathbf{r}^{\prime})\rangle$ represents the correlator generated from the connected generating functional, which removes a coordinate-independent constant. This formula can be employed to evaluate the coherence length for several lattice model, such as sawtooth and Lieb lattices.
Projected mean-field Hamiltonian and Green function
@@
$a_{i\alpha\sigma}^{\dagger}\rightarrow\frac{1}{\sqrt{N}}\sum_{\bm{k}}e^{-i(\bm{r}_{i}+\bm{\delta}_{\alpha})\cdot\bm{k}}u_{\alpha\sigma}(\bm{k})c_{\sigma}^{\dagger}(\bm{k}),$ (14)
@@
$H_{0}=\sum_{\bm{k}\sigma}\epsilon(\bm{k})c_{\bm{k}\sigma}^{\dagger}c_{\bm{k}\sigma}+\sum_{\bm{k}}\left[c_{\uparrow}^{\dagger}(\bm{k})c_{\downarrow}^{\dagger}(-\bm{k})\Delta_{0}+h.c.\right],$ (15)
@@
$\Delta_{0}=\frac{1}{N}\sum_{\alpha}\sum_{\bm{k}}\Delta_{\alpha}u_{\alpha}(\bm{k})u_{\alpha}^{*}(\bm{k}).$ (16)
@@
$\Delta=-\frac{U}{N}\sum_{\bm{k}}\langle c_{\downarrow}(-\bm{k})c_{\uparrow}(\bm{k})\rangle,$ (17)
@@
$=\sum_{\bm{k}}\psi_{\bm{k}}^{\dagger}[(\epsilon(\bm{k})-\mu)\tau_{z}+\Delta\tau_{x}]\psi_{\bm{k}}+\sum_{\bm{k}}(\epsilon(\bm{k})-\mu)+\frac{1}{U}\Delta^{2},$ (18)
@@
$G(\bm{r}\tau;\bm{r}^{\prime}\tau^{\prime})=\begin{bmatrix}\langle\hat{T}[c_{\uparrow}(\bm{r},\tau)c_{\uparrow}^{\dagger}(\bm{r}^{\prime},\tau^{\prime})]\rangle&\langle\hat{T}[c_{\uparrow}(\bm{r},\tau)c_{\downarrow}(\bm{r}^{\prime},\tau^{\prime})]\rangle\\ \langle\hat{T}[c_{\downarrow}^{\dagger}(\bm{r},\tau)c_{\uparrow}^{\dagger}(\bm{r}^{\prime},\tau^{\prime})]\rangle&\langle\hat{T}[c_{\downarrow}^{\dagger}(\bm{r},\tau)c_{\downarrow}(\bm{r}^{\prime},\tau^{\prime})]\rangle\end{bmatrix},$ (19)
@@
$G(i\omega_{n},\bm{k})=\frac{-i\omega_{n}\tau_{0}-(\epsilon(\bm{k})-\mu)\tau_{z}-\Delta\tau_{x}}{\omega_{n}^{2}+(\epsilon(\bm{k})-\mu)^{2}+\Delta^{2}},$ (20)
where $\omega_{n}=(2n+1)\pi T$ is the Matsubara frequency and $\tau_{0}$ is a $2\times 2$ identity matrix. With the Gor’kov’s Green function in Eq. (20), we can determine the order parameters by
$\Delta=-\frac{U}{2N}T\sum_{\bm{k},n}\mathrm{Tr}G(i\omega_{n},\bm{k})\tau_{x}=\frac{U}{N}\sum_{\bm{k},n}\frac{\Delta}{\omega_{n}^{2}+(\epsilon(\bm{k})-\mu)^{2}+\Delta^{2}},$ (21)
@@
$1=\frac{U}{N}T\sum_{\bm{k},n}\frac{1}{\omega_{n}^{2}+(\epsilon(\bm{k})-\mu)^{2}+\Delta^{2}}.$ (22)
@@
$2\nu=1+\frac{T}{N}\sum_{\bm{k},n}\mathrm{Tr}G(i\omega_{n},\bm{k})\tau_{z}.$ (23)
with $\nu$ being the filling factor. Combine Eqs. (22) and (23) and we can obtain the superconducting order parameter.
Interaction renormalized chemical potential $\mu$
@@
$\nu$ $=\frac{1}{2}\left(1-\text{$\int\frac{d^{2}k}{(2\pi)^{2}}$}\frac{\epsilon(\bm{k})-\mu_{0}}{|\epsilon(\bm{k})-\mu_{0}|}\tanh\frac{\beta|\epsilon(\bm{k})-\mu_{0}|}{2}\right),$ (24)
$\nu$ $=\frac{1}{2}\left(1-\int\frac{d^{2}k}{(2\pi)^{2}}\frac{\epsilon(\bm{k})-\mu}{\varepsilon(\bm{k})}\tanh\frac{\beta\varepsilon(\bm{k})}{2}\right).$ (25)
@@
$=\frac{m}{2\pi}\left(\sqrt{\Delta^{2}}-\sqrt{\Delta^{2}+\mu^{2}}\right).$ (26)
@@
$\mu_{0}=\Delta-\sqrt{\Delta^{2}+\mu^{2}}=\frac{\mu^{2}}{\Delta+\sqrt{\Delta^{2}+\mu^{2}}},$ (27)
@@
Supplementary Note 2: Quantum geometry and coherence length
Proof of the independence of coherence length from orbital positions